# TinyIO: Lightweight Reparameterized Inertial Odometry

### v2

@@ 3.2 Comparisons with the baselines
Metric evaluation. Table 1 (left) summarizes ATE and RTE comparisons between TinyIO and the baseline models. TinyIO and its reparameterized variant, TinyIO_deploy, achieve the lowest localization errors in most scenarios. For example, relative to RoNIN ResNet, TinyIO reduces ATE by 2.59%—13.54% and RTE by 4.82%—9.33% across the four datasets. Model complexity increases progressively from RoNIN to TLIO and RNIN, which raises deployment challenges for edge devices. Although IMUNet uses depthwise-separable convolutions to reduce complexity, its localization accuracy remains limited. In contrast, TinyIO decouples training and inference to substantially reduce parameter counts, while the SGU suppresses implicit noise and fuses local fine-grained and contextual motion features to strike an effective balance between accuracy and compactness. Overall, these improvements explain the superior performance of TinyIO over the baselines.
@@
Visualization. Fig. 5 shows a comparison of PDE among TinyIO, TinyIO_deploy, and RoNIN (results on the RoNIN dataset are shown due to space constraints). TinyIO attains the smallest PDE median and the most compact error distribution, indicating superior endpoint approximation. Although reparameterization slightly increases PDE for TinyIO_deploy, TinyIO_deploy still outperforms RoNIN ResNet and preserves high-fidelity trajectory reconstruction consistent with the training-phase model.
@@
Fig. 4 visualizes trajectory reconstructions produced by TinyIO and the baseline models on the test sets. In Fig. 4(d), TinyIO effectively captures contextual motion information, yielding accurate and stable tracking. After reparameterization, the positioning error of TinyIO_deploy increases slightly but remains lower than that of RoNIN ResNet, suggesting that the accuracy degradation due to reparameterization is within acceptable bounds. For complex trajectories (Fig. 4(a)–(c)), TinyIO better models local fine-grained motion and exhibits substantially less drift than RoNIN ResNet when handling rotational or highly nonlinear motions; consequently, both TinyIO and TinyIO_deploy closely follow ground-truth trajectories, whereas RoNIN ResNet shows pronounced drift during turns and increasingly larger deviations as trajectory complexity grows.
3.3 Ablation study
@@
We conduct ablation experiments to validate the reparameterization strategy and the contributions of individual modules (SGU and RepBlock).
@@
The middle section of Table 1 reports performance comparisons between TinyIO and TinyIO_deploy across the four datasets. After reparameterization, TinyIO_deploy (3.809M parameters) reduces the parameter count by approximately 14.52% relative to TinyIO (4.456M), and its model size is only 82.18% that of RoNIN ResNet (4.635M). Despite this compression, the increases in ATE and RTE for TinyIO_deploy are marginal, and TinyIO_deploy still outperforms RoNIN ResNet in localization accuracy. Visualization results in Fig. 4 corroborate these findings. These results indicate that the proposed reparameterization achieves a favorable trade-off between parameters and accuracy, making TinyIO particularly suitable for edge devices with limited storage.
@@
The right portion of Table 1 shows module ablation results. Removing SGU (TinyIO w/o SGU) leaves only the RepBlock: the model still captures local fine-grained motion and yields substantially lower localization error than RoNIN ResNet, indicating that RepBlock both reduces parameter complexity and contributes positively to accuracy. Integrating SGU into the full TinyIO architecture further improves accuracy, demonstrating that multi-scale feature fusion via SGU is critical for enhanced trajectory estimation. These findings confirm the effectiveness and complementarity of the proposed components.
4 Conclusion
@@
5 Acknowledgements
This work is supported by Science and Technology Major Program of Fujian Province (No. 2022HZ026007) and partly supported by the Science and Technology Planning Project of Fujian province (2022I0001).

### v4

@@ IV-B Comparisons with the baselines
Metric evaluation. Table I summarizes ATE and RTE comparisons between TinyIO and the baseline models. TinyIO and its reparameterized variant, TinyIO_deploy, achieve the lowest localization errors in most scenarios. For example, relative to R-ResNet, TinyIO reduces ATE by 22.22%—59.71% and RTE by 3.68%—28.46% across the all datasets. Model complexity increases progressively from RoNIN to TLIO and RNIN, which raises deployment challenges for edge devices. Although IMUNet uses depthwise-separable convolutions to reduce complexity, its localization accuracy remains limited. In contrast, TinyIO decouples training and inference to substantially reduce parameter counts, while the DPAA models contextual motion information through a lightweight attention mechanism to strike an effective balance between accuracy and compactness. Overall, these improvements explain the superior performance of TinyIO over the baselines.
@@
Visualization. Fig. 4(a) and (b) present the cumulative distribution function (CDF) curves of the ATE and RTE for TinyIO, TinyIO_deploy, R-LSTM and R-ResNet on the IMUNet dataset, respectively. Results on the IMUNet dataset are presented due to space constraints. As shown in the figures, the TinyIO curve (orange) is positioned above and to the left of the R-ResNet curve (blue) and R-LSTM (green), indicating that TinyIO achieves higher accuracy compared to R-ResNet and R-LSTM. For instance, approximately 82% of trajectories from TinyIO exhibit ATE values below 0.06 m, whereas only approximately 72% of trajectories from R-ResNet achieve ATE values below this threshold. Fig. 4(c) shows a comparison of PDE. TinyIO attains the smallest PDE median and the most compact error distribution, indicating superior endpoint approximation. Although reparameterization slightly increases the PDE for TinyIO_deploy, it still outperforms R-ResNet and preserves high-fidelity trajectory reconstruction consistent with the training-phase model.
TABLE II: Ablation study results on open-source datasets. Best results are highlighted in orange, second-best in green.
Dataset Metric R-ResNet TinyIO (only RepBlock) Full TinyIO (RepBlock&DPAA) Performance (meters) RoNIN ATE 6.12 4.99 4.68 RTE 4.08 3.94 3.93 RIDI ATE 2.43 2.16 1.89 RTE 2.60 2.50 2.19 IMUNet ATE 8.16 7.45 5.66 RTE 4.71 5.13 4.35 OxIOD ATE 3.50 2.28 1.41 RTE 1.30 1.12 0.93 Parameters (M) 4.635 3.533 3.57 FLOPs (M) 38.252 26.019 26.194 Peak Memory (MB) 29.069 25.587 25.609
@@
Fig. 5 visualizes trajectory reconstructions produced by TinyIO and the baseline models on the test sets. In Fig. 5(a), TinyIO effectively captures contextual motion information, yielding accurate and stable tracking. After reparameterization, the positioning error of TinyIO_deploy increases slightly but remains lower than that of R-ResNet, suggesting that the accuracy degradation due to reparameterization is within acceptable bounds. For complex trajectories (Fig. 5(b), (c)), TinyIO better models local fine-grained motion and exhibits substantially less drift than R-ResNet when handling rotational or highly nonlinear motions; consequently, both TinyIO and TinyIO_deploy closely follow ground-truth trajectories, whereas R-ResNet shows pronounced drift during turns and increasingly larger deviations as trajectory complexity grows.
IV-C Ablation study
@@
We conduct ablation experiments to validate the reparameterization strategy and the contributions of individual modules (DPAA and RepBlock).
@@
The Table I reports performance comparisons between TinyIO and TinyIO_deploy across the all datasets. After reparameterization, TinyIO_deploy (2.975M parameters) reduces the parameter count by approximately 16.67% relative to TinyIO (3.570M), and its model size is only 64.19% that of R-ResNet (4.635M). Despite this compression, the increases in ATE and RTE for TinyIO_deploy are marginal, and TinyIO_deploy still outperforms R-ResNet in localization accuracy. Visualization results in Fig. 5 corroborate these findings. These results indicate that the proposed reparameterization achieves a favorable trade-off between parameters and accuracy, making TinyIO particularly suitable for edge devices with limited storage.
@@
The Table II shows module ablation results. Removing DPAA (TinyIO (only RepBlock)) leaves only the RepBlock: the model still captures local fine-grained motion and yields substantially lower localization error than R-ResNet, indicating that RepBlock both reduces parameter complexity and contributes positively to accuracy. Integrating DPAA into the full TinyIO (RepBlock&DPAA) further improves accuracy, indicating that modeling contextual motion through DPAA is essential for enhancing trajectory estimation. These findings confirm the effectiveness and complementarity of the proposed components.
V Conclusion